Find the values of m for which mx^2 − 6x + 5 = 0 has one solution.
Answer by htmentor(1343)

There will be one solution if the discriminant b^2 - 4ac = 0.
(-6)^2 - (4)(5)m = 0
36 = 20m
m = 36/20 = 9/5